OpenStudy (anonymous):

graph the equation by plotting 3 points -3y=-2x-6

OpenStudy (anonymous):

Well, first you find the x and y intercept (which adds up to 2 points so far) then you put in a value in x and get y. (your final point) -3y=-2x-6 y=2/3x+2 y intercept: 0,2 x intercept: 0=2/3x+2 -2=2/3x =3=x